The diameter of Cylinder I is 5 and its height is 9.The diameter of Cylinder II is
10 and its height is 9.What is the ratio of the volume of Cylinder I to Cylinder II?
(A) 1 : 2 (B) 1 : 4 (C) 1 : 5 (D) 1 : 8 (E) 1 : 25

Answer by mananth(16946)

The diameter of Cylinder I is 5 and its height is 9.
The diameter of Cylinder II is 10 and its height is 9.
What is the ratio of the volume of Cylinder I to Cylinder II?
..
Volume of cylinder I = pi*(2.5)^2*9
Volume of cylinder II = pi*(5)^2*9
Ratio = pi*(2.5)^2*9/ pi*(5)^2*9
Ratio = 6.25/25
Ratio =1/4